OpenStudy (anonymous):

Solve the equation for the letter y: -2x + 4y = 8 y = 2x + 8 y = 4(2x + 8) y = -2x - 8 none of the above

OpenStudy (anonymous):

-2x + 4y = 8 -2x -2x 4y = -2x + 8 divide both sides by 4 y = -1/2x + 2 is your answer

OpenStudy (anonymous):

none of the above

OpenStudy (promisespromises):

Actually you would add the 2x so -2x + 4y = 8 +2x +2x 4y=2x+8 y=2/4x +8/4 y=1/2x+2 According to your choices it's none of the above.